OpenStudy (anonymous):

Solve. Check for extraneous solutions. Did I do this problem correctly? 1. (x+5)^1/2 - (5-2x)^1/4 = 0 ((x+5)^1/2)^2 - ((5-2x)^1/4)^4 = 0 (x+5) - (5-2x) = 0 x+5 = 5 +2x 5 = 5+x x = 0 I'm not sure if I did my math correctly.

OpenStudy (anonymous):

nope

OpenStudy (anonymous):

Oop I thought so. It didn't seem right. Do you mind explaining?

OpenStudy (anonymous):

you cant multiply powers like that

OpenStudy (anonymous):

hold on i will tell

OpenStudy (anonymous):

Thank you (:

OpenStudy (anonymous):

\[(x+5)^{1/2} = (5-2x)^{1/4}\]

OpenStudy (anonymous):

see i tok 5-2x term on right side

OpenStudy (anonymous):

took*

OpenStudy (anonymous):

now

OpenStudy (anonymous):

multiply the powers by 4 \[(x+5)^{4/2}=(5-2x)^{4/4}\]

OpenStudy (anonymous):

you get \[(x+5)^2 = 5-2x\] solve the rest tell me the answer

OpenStudy (anonymous):

understood?

OpenStudy (anonymous):

Oh yes I see! Thank you very much! (:

OpenStudy (anonymous):

so what are the answers?

OpenStudy (anonymous):

I'm working on it. So far I have x^2 +10+25=5-2x

OpenStudy (anonymous):

okay let me know when you are done

OpenStudy (anonymous):

Okay I am done. I did: x^2 + 12x + 20 = 0 (x+10)(x+2) = 0 x= -10 x=-2 BUT -10 is an extraneous solution because if you substitute it for x, it won't equal to 0.

OpenStudy (anonymous):

why wouldn't it? \[-10^2 +12(-10)+20 =100-120+20 =120-120=0\]

OpenStudy (anonymous):

right?

OpenStudy (anonymous):

I substituted it back to the original problem (x+5)^1/2 - (5-2x)^1/4 = 0 and the calculator just said domain error.

OpenStudy (anonymous):

LOL why to put it in original equation when you have simplified 1 -10 and -2 are your values

OpenStudy (anonymous):

Yes that is correct. My paper just asked me to check for extraneous solutions and that's what I've been doing for my other problems. Thank you for the big help though (: I was so lost :o

OpenStudy (anonymous):

in original equation -10+5 will give you -5 so sqrt of -5 is imaginary number that's why it gives domain error

OpenStudy (anonymous):

haha you are welcome x)